OpenStudy (anonymous):

Sophie works 4 hours a day, and earns a salary of $500.00 per month. She earns $10 for every extra hour that she works. If she earned $650.00 last month, how many extra hours did Sophie work?

OpenStudy (anonymous):

@ganeshie8 is the answer 15?

OpenStudy (nikato):

yea, i got that too

ganeshie8 (ganeshie8):

extra money earned = 650-500 = $150 since each hour she gets $10, hours worked = 150/10 = 15 hours

ganeshie8 (ganeshie8):

is that how you arrived at 15 ?

OpenStudy (anonymous):

yes! so is it correct?

ganeshie8 (ganeshie8):

Yep! good work !